My Cart
Your Cart 0

    Your cart is empty.

  • Total (Amount) ₹0.00
NCERT · Class 11 · Maths

Chapter 14: Probability

Practice each question, then open the answer or worked solution.

Exercise 14.2

21 questions
Q11Written answer

In a lottery, a person choses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? (Hint: order of the numbers is not important.)

Show solutionHide solution
Worked solution

Ans: According to question.

Total numbers of numbers in the draw is 20

Numbers to be selected is 6

Therefore,

\[n\left( S \right){ = ^{20}}{C_6}\]

\[n\left( S \right)=\frac{\left| \underline{20} \right.}{\left| \underline{6\left| \underline{20-6} \right.} \right.}\]

\[n\left( S \right) = 38760\]

Let  be the event that six numbers match with the six numbers already fixed by the lottery committee.

\[n\left( A \right){ = ^6}{C_6}\]

\[n\left( A \right) = 1\]

Therefore, the probability of winning the prize,

\[P\left( A \right) = \dfrac{{n\left( A \right)}}{{n\left( S \right)}}\]

\[P\left( A \right) = \dfrac{1}{{38760}}\]

 

Q12Written answer

Check whether the following probabilities P(A) and P(B) are consistently defined

(i) \[P\left( A \right) = 0.5,{\text{ }}P\left( B \right) = 0.7,{\text{ }}P\left( {A \cap B} \right) = 0.6\]

(ii) \[P\left( A \right) = 0.5,{\text{ }}P\left( B \right) = 0.4,{\text{ }}P\left( {A \cup B} \right) = 0.8\]

Show solutionHide solution
Worked solution

Ans:\[P\left( {A \cap B} \right) > P\left( A \right)\]

Therefore, the given probabilities are not consistently defined.

Ans: As we know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[0.8 = 0.5 + 0.4--P(A \cap B)\]

\[0.8 - 0.9 = --P(A \cap B)\]

\[P(A \cap B) = 0.1\]

Therefore, \[P\left( {A \cap B} \right) < P\left( A \right)\] and \[P\left( {A \cap B} \right) < P\left( B \right)\]

So, the given probabilities are consistently defined.

 

Q13Written answer

Fill in the blanks in following table:


P(A)

P(B)

P(A ⋂ B)

P(A ⋃ B)

(i)

1/3

1/5

1/15

….

(ii)

0.35

…..

0.25

0.6

(iii)

0.5

0.35

….

0.7

 

Show solutionHide solution
Worked solution

Ans:

(i) We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[P(A \cup B) = \dfrac{1}{3} + \dfrac{1}{5} + \dfrac{1}{{15}}\]

\[P(A \cup B) = \dfrac{7}{{15}}\]

Ans:

(ii) We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[0.6 = 0.35 + P\left( B \right) - 0.25\]

\[P\left( B \right) = 0.6 - 0.35 + 0.25\]

\[P\left( B \right) = 0.5\]

Ans: 

(iii) We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[0.7 = 0.5 + 0.35 - P(A \cap B)\]

\[P(A \cap B) = 0.5 + 0.35 - 0.7\]

\[P(A \cap B) = 0.15\]

 

Q14Written answer

Given P(A) = 5/3 and P(B) = 1/5 . Find P(A or B), if A and B are mutually exclusive events.

Show solutionHide solution
Worked solution

Ans: \[P\left( A \right) = \dfrac{3}{5}\] and \[P\left( B \right) = \dfrac{1}{5}\]

If A and B are mutually exclusive

\[P(A \cup B) = P\left( A \right) + P\left( B \right)\]

\[P(A \cup B) = \dfrac{3}{5} + \dfrac{1}{5}\]

\[P(A \cup B) = \dfrac{4}{5}\]

 

Q15Written answer

If E and F are events such that P(E) = ¼ , P(F) = ½ and P(E and F) = 1/8 , find

(i) P(E or F)

(ii) P(not E and not F)

Show solutionHide solution
Worked solution

Ans: We have, \[P\left( E \right) = \dfrac{1}{4}\], \[P\left( F \right) = \dfrac{1}{2}\], \[P\left( {E \cap F} \right) = \dfrac{1}{8}\]

We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[P(A \cup B) = \dfrac{1}{4} + \dfrac{1}{2} - \dfrac{1}{8}\]

\[P(A \cup B) = \dfrac{5}{8}\]

Ans: \[P\left( {{\text{not E and not F}}} \right) = P\left( {\overline E  \cap \overline F } \right)\]

\[P\left( {{\text{not E and not F}}} \right) = P\left( {\overline {E \cup F} } \right)\]

\[P\left( {{\text{not E and not F}}} \right) = 1 - P\left( {E \cup F} \right)\]

\[P\left( {{\text{not E and not F}}} \right) = 1 - \dfrac{5}{8}\]

\[P\left( {{\text{not E and not F}}} \right) = \dfrac{3}{8}\]

 

Q16Written answer

Events E and F are such that P(not E or not F) = 0.25, State whether E and F are mutually exclusive.

Show solutionHide solution
Worked solution

Ans: We have,

\[P\left( {{\text{not E or not F}}} \right) = 0.25\]

\[P\left( {\overline {\text{E}}  \cup \overline {\text{F}} } \right) = 0.25\]

\[P\left( {\overline {E \cap F} } \right) = 0.25\]

\[1 - P\left( {E \cap F} \right) = 0.25\]

\[P\left( {E \cap F} \right) = 1 - 0.25\]

\[P\left( {E \cap F} \right) = 0.75\]

\[P\left( {E \cap F} \right) \ne 0\]

Therefore, E and F are not mutually exclusive event.

 

Q17Written answer

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine

(i) P(not A)

(ii) P(not B)

(iii) P(A or B)

Show solutionHide solution
Worked solution

Ans: We know that,

\[P\left( {{\text{not A}}} \right) = 1 - P\left( A \right)\]

\[P\left( {{\text{not A}}} \right) = 1 - 0.42\]

\[P\left( {{\text{not A}}} \right) = 0.58\]

Ans: We know that,

\[P\left( {{\text{not B}}} \right) = 1 - P\left( B \right)\]

\[P\left( {{\text{not B}}} \right) = 1 - 0.48\]

\[P\left( {{\text{not B}}} \right) = 0.52\]

Ans: We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[P(A \cup B) = 0.42 + 0.48 - 0.16\]

\[P(A \cup B) = 0.74\]

 

Q18Written answer

In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.

Show solutionHide solution
Worked solution

Ans: Let A be the event that the student is studying mathematics and B be the event that the student is studying biology,

Therefore,

\[P\left( A \right) = \dfrac{{40}}{{100}}\]

\[P\left( A \right) = \dfrac{2}{5}\]

And,

\[P\left( B \right) = \dfrac{{30}}{{100}}\]

\[P\left( B \right) = \dfrac{3}{{10}}\]

So, student studying both subjects,

\[P\left( {A \cap B} \right) = \dfrac{{10}}{{100}}\]

\[P\left( {A \cap B} \right) = \dfrac{1}{{10}}\]

Therefore, probability of studying mathematics or biology is given as,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[P(A \cup B) = \dfrac{2}{5} + \dfrac{3}{{10}} - \dfrac{1}{{10}}\]

\[P(A \cup B) = \dfrac{6}{{10}}\]

\[P(A \cup B) = \dfrac{3}{5}\]

Therefore, \[\dfrac{3}{5}\] is the probability that student will studying mathematics or biology.

 

Q19Written answer

In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?

Show solutionHide solution
Worked solution

Ans: Let probability of a randomly chosen student passing the first examination is 0.8 be \[P\left( A \right)\].

And also probability of passing the second examination is 0.7 be \[P\left( B \right)\]

Therefore,

\[P(A \cup B)\;\]is probability of passing at least one of the examination

So, We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[0.95 = 0.8 + 0.7 - P(A \cap B)\]

\[P(A \cap B) = 0.8 + 0.7 - 0.95\]

\[P(A \cap B) = 0.55\]

therefor, 0.55 is the probability that student will pass both the examinations.

 

Q20Written answer

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?

Show solutionHide solution
Worked solution

Ans: Let probability of passing the English examination is 0.75 be \[P\left( A \right)\].

And also assume the probability of passing the Hindi examination is be \[P\left( B \right)\].

We hve given, \[P\left( A \right) = 0.75,{\text{ }}P\left( {A \cap B} \right) = 0.5,{\text{ }}P\left( {A' \cap B'} \right) = 0.1\]

We know that, 

\[P\left( {A' \cap B'} \right) = 1--P(A \cup B)\]

\[P(A \cup B) = 1--P\left( {A' \cap B'} \right)\]

\[P(A \cup B) = 1 - 0.1\]

\[P(A \cup B) = 0.9\]

We know that,

\[P(A \cup B) = P\left( A \right) + P\left( B \right)--P(A \cap B)\]

\[0.9 = 0.75 + P\left( B \right) - 0.5\]

\[P\left( B \right) = 0.9 + 0.5 - 0.75\]

\[P\left( B \right) = 0.65\]

Therefore, the probability of passing  the hindi examination is 0.65.